OpenStudy (anonymous):

a coin has the numbers 0.4 printed on one side and 0.9 printed on the other side. Suppose that the coin is weighted so that when it is tossed, the side with the number .4 appears 80% of the time

OpenStudy (anonymous):

if the coin is tossed once, find the probability that the number shown will be greater than .55

OpenStudy (anonymous):

if the coin is tossed 25 times, estimate the probability that the average of the 25 numbers shown will be greater than .55

OpenStudy (kropot72):

When the coin is tossed there are only 2 possible outcomes: 0.4 or 0.9. P(0.4) = 0.8 P(0.9) = 1.0 - P(0.4)

OpenStudy (anonymous):

so .2

OpenStudy (kropot72):

P(0.9) = 0.2 is correct for the fist part.

OpenStudy (anonymous):

ok cool

OpenStudy (anonymous):

idk how to do the second part

OpenStudy (kropot72):

In 25 tosses the estimated sum of the 0.4's is 25 * 0.8 * 0.4 In 25 tosses the estimated sum of the 0.9's is 25 * 0.2 * 0.9 The estimated average of the 25 numbers shown is given by \[\frac{(25\times 0.8\times 0.4)+(25\times 0.2\times 0.9)}{25}=you\ can\ calculate\]